OpenStudy (anonymous):

Factor completely: -4x2 + 16x - 24

OpenStudy (ghostgate):

I may be incorrect by the way. Its been awhile. -4 x 2 = -8 + 16x - 24 Well 24 + 8 = 32. 32 / 16 = 2. So if I'm right x = 2. Hope that helps. :)